# Expert-Token Resonance: Redefining MoE Routing through Affinity-Driven Active Selection

### v2

@@ Abstract
Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency. However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization. We introduce a novel framework that redefines MoE routing through affinity-driven active selection. The innovations for the framework encompass: (1) A rigorous formulation of expert-token affinity metrics. (2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens. (3) Theoretical derivation and experimental evidence of reduced expert capacity bounds under dynamic token distribution evolution. It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization. Our theoretical analysis demonstrates that this approach mitigates expert homogenization while enabling substantial capacity boundary reduction. Experimental validation corroborates these findings: it achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy. When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed. After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, C-Eval, and TeleQnA benchmarks.
@@ Introduction
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al. 2023). As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al. 2024b). To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al. 2020). The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022). For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al. 2022). Recently, several LLMs employing MoE structures, such as DeepSeek-MoE (Dai et al. 2024) and Mixtral (Jiang et al. 2024a), have demonstrated outstanding performance on various leaderboards.
@@
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al. 2017). The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts. The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies. Recent studies address these challenges from multiple perspectives (Li et al. 2023). DeepSeek-MoE (Dai et al. 2024) introduces the concept of the vertical-class expert, which is created through fine-grained partitioning, and the shared experts are always activated. This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy. StableMoE (Dai et al. 2022) proposes a two-stage training approach to address the issue of routing fluctuation. This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens. The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts. To address this, MoDE (Xie et al. 2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths. DYNMoE (Guo et al. 2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
@@
In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC). Google Brain introduces the EC routing algorithm (Zhou et al. 2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. The Brainformer (Zhou et al. 2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert. Then, each token is routed to the Top-k experts. This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes.
@@
The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al. 2021). In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model. Early work introduced the concept of expert capacity (Lepikhin et al. ), which refers to the maximum number of tokens each expert can process at once. Tokens exceeding this capacity are discarded. Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation. However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes. XMoE (Yang et al. 2024) achieves more precise router by implementing a threshold-based approach. If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection. This method allows for more nuanced token selection and processing. LocMoE (Li et al. 2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction. Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
@@
    A rigorous formulation of expert-token affinity metrics. Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts. We provide theoretical analysis of utilizing the affinity score for better guiding expert choice routers to focus on different segmentations of tokens, thus mitigating the homogenization tendency. Further, we incorporate affinity scores into the design of our bidirectional selection router.
@@
Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters. Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section Related Work introduces related work on LLMs, MoE, and the Ascend architecture. Section Method reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence. Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
Related Work
LLMs.
BERT (Kenton and Toutanova 2019) and GPT (Radford et al. ) in 2018 marked the emergence of the LLM era. GPT-3 (Floridi and Chiriatti 2020), with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020. By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 (Achiam et al. 2023) achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 (Anil et al. 2023) enhanced knowledge acquisition and reasoning; Chinchilla (Hoffmann et al. 2022) optimized computational resources and model scale; and Claude-v1 (Wei, Haghtalab, and Steinhardt 2024) displayed a safer, human-aligned LLM. These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
MoE.
Despite the outstanding performance of LLMs in language understanding and generation, their high computational cost hinders further model scale expansion. To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance. GShard (Lepikhin et al. 2020) and Switch Transformer (Fedus, Zoph, and Shazeer 2022) employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem. Hash Layer (Roller et al. 2021) adopt hashing techniques to simplify the gating mechanism and improve model performance. METRO (Bajaj et al. 2022) introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability. DSelect-k (Hazimeh et al. 2021) proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs. Furthermore, DeepSpeed-MoE (Rajbhandari et al. 2022) optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency. These works lay the foundation for the subsequent application of MoE in LLMs.
Ascend Architecture.
The AI Core in the Ascend 910B NPU adopts the DaVinci architecture (Tang and Wang 2023), which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training. Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) (Xia et al. 2021), enables high-performance inter-device data communication in multi-card scenarios. The Huawei Collective Communication Library (HCCL) (Cao et al. 2024), designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication. It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training. PyTorch (Paszke et al. 2019), a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch (Zhu et al. 2023). AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication. ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
@@
In this section, we present the definition of affinity score with a rigorous formulation of expert-token affinity metrics and the implementation of our affinity-driven active selection routing. To reduce the lower bound of expert capacity, we introduce a training strategy for dynamic token distribution evolution under the guiding principle of affinity score.
@@
where $n$ is the number of experts, $G(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}^{n}$ is the gating weight vector of experts which maps the tokens of $\bm{x}_{t}$ into the coresponding experts with weights, e.g., $G_{i}(\bm{x})=\text{Softmax}(\bm{W}\bm{x}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}$ is the $i$-th expert network, see (Liu et al. 2024) for current different router methods. Generally, $n\ll s$, which saves much computation compared to the dense structure.
Expert-Token Affinity.
$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix. With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al. 2017) has been proposed. Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
@@
The Dynamic Token Distribution Evolution.
Under the premises of orthogonal gating weights and a data distribution approaching uniformity, LocMoE demonstrates that the expert capacity is closely related to the angle between the gating weights and tokens. For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
@@
This study employs the Mixtral 8×7B model, incorporating LocMoE and our proposed approach. The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing. Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128. Other details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix.
@@
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval (Huang et al. 2024) evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA (Maatouk et al. 2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
@@
GDAD-1 GDAD-2 GDAD-3 Avg C-Eval TeleQnA
Baseline 47.8 43.0 65.4 52.8 38.5 62.1
LocMoE 55.5 47.6 71.1 59.0 42.6 67.6
LocMoE+ 57.4 49.9 74.5 61.5 43.1 68.8

### v3

# Expert-Token Resonance MoE: Bidirectional Routing with Efficiency Affinity-Driven Active Selection

@@ Abstract
Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency. However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization. We propose a novel expert routing framework that incorporates: (1) An efficient routing mechanism with lightweight computation. (2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens. (3) A module that determines the lower bounds of expert capacity based on dynamic token distribution analysis, specifically designed to address drop-and-pad strategies. It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization. This framework effectively reduces expert homogeneity while enhancing the performance of the expert selection module. Additionally, we introduce a local expert strategy that simultaneously improves load balancing and reduces network communication overhead. It achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy. When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed. After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, GPQA, and TeleQnA benchmarks.
@@ Introduction
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al. 2023). As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al. 2024b). To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al. 2020). The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022). For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al. 2022). Recently, several LLMs employing MoE structures, such as DeepSeek-V3 (Liu et al. 2024a) and Mixtral (Jiang et al. 2024a), have demonstrated outstanding performance on various leaderboards.
@@
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al. 2017). The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts. The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies. Recent studies address these challenges from multiple perspectives (Li et al. 2023). StableMoE (Dai et al. 2022) proposes a two-stage training approach to address the issue of routing fluctuation. This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens. Dynamic-MoE (Huang et al. 2024a) designs a dynamic routing Mixture-of-Experts (MoE) policy that evaluates the sufficiency of current experts while reducing activated parameters by 90%. The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts. To address this, MoDE (Xie et al. 2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths. DYNMoE (Guo et al. 2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
@@
In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC). Google Brain introduces the EC routing algorithm (Zhou et al. 2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. The Brainformer (Zhou et al. 2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert. Then, each token is routed to the Top-k experts. This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes. Autonomy-of-Experts models (Lv et al. 2025) design a novel MoE paradigm in which experts autonomously select themselves to process inputs by aware of its own capacity to effectively process a token.
@@
The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al. 2021). In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model. Early work introduced the concept of expert capacity (Lepikhin et al. ), which refers to the maximum number of tokens each expert can process at once. Tokens exceeding this capacity are discarded. Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation. However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes. DeepSeek-V2 (Liu et al. 2024a) designs a device-limited routing mechanism to bound MoE-related communication cost. DeepSeek-V3 (Liu et al. 2024a) pioneers an auxiliary-loss-free strategy to minimizes the performance degradation. This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy. XMoE (Yang et al. 2024) achieves more precise router by implementing a threshold-based approach. If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection. This method allows for more nuanced token selection and processing. LocMoE (Li et al. 2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction. Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
@@
    Affinity-based Efficient Expert Routing via GrAP. By leveraging cosine similarity between tokens and gating weights to define affinity scores, our router effectively guides experts to focus on distinct token segments, mitigating the expert homogenization problem. Meanwhile, the GrAP design reduces computational complexity by a factor of 1/2D to 1/D compared to traditional MLPs (D denotes the dimension of the intermediate hidden layer). This integrated approach demonstrates both improved routing effectiveness and substantial computational efficiency.
@@
Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes MindSpeed-LLM, MindSpeed, and Megatron-LM (Shoeybi et al. 2019) libraries for training on Ascend NPU clusters. Ascend designs a new computing architecture for LLM training and inference scenarios (Liao et al. 2021), boasting powerful low-bit computing capabilities. Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section Method presents the methods proposed in this paper, along with theoretical evidence. Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
@@
In this section, we present the efficient routing mechanism, and our adaptive bidirectional selection mechanism is detailed. Then, for traditional drop-and-pad strategies, a dynamic token distribution analysis module that optimizes the lower bounds of expert capacity are displayed. Moreover, we also describe the loss for expert load balancing.
@@
where $n$ is the number of experts, $G(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}^{n}$ is the gating weight vector of experts which maps the tokens of $\bm{x}_{t}$ into the coresponding experts with weights, e.g., $G_{i}(\bm{x})=\text{Softmax}(\bm{W}\bm{x}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}$ is the $i$-th expert network, see (Liu et al. 2024b) for current different router methods. Generally, $n\ll s$, which saves much computation compared to the dense structure.
Cost-Efficient Sparse Expert-Token Affinity.
$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix. After processing through the GrAP routing layer, tokens generate a diagonal sparse matrix as shown. Compared to the dense matrix produced by traditional routing layers, this reduces the parameter count to 1/D of the original, significantly decreasing the computational overhead of the expert routing layer.
With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al. 2017) has been proposed. Building upon the auxiliary loss, our work introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
@@
Token Distribution Dynamics under Expert Routing.
Under the premises of orthogonal gating weights and a data distribution approaching uniformity, the previous studies demonstrate that the expert capacity is closely related to the angle between the gating weights and tokens. For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
@@
This study employs the Mixtral 8×7B model, incorporating our proposed approach. The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing. Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. Other details of experimental setup including datasets, environment, and metrics, can be seen in Appendix.
@@
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 13.6% compared to the Baseline and 2.8% compared to LocMoE in the GPQA (Rein et al. 2023) evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA (Maatouk et al. 2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
@@
GDAD-1 GDAD-2 GDAD-3 Avg GPQA TeleQnA
Baseline 47.8 43.0 65.4 52.8 29.5 62.1
LocMoE 55.5 47.6 71.1 59.0 32.6 67.6
LocMoE+ 57.4 49.9 74.5 61.5 33.5 68.8